Borland, Inc., has a profit margin of 6.75 percent on sales of $23,600,000. Assume the firm has debt of $9,200,000 and total assets of $15,800,000.

What is the firm’s ROA?
Answer #1

Profit margin=net income/sales

net income=23,600,000*6.75%=$1593000

ROA=net income/Total assets

=1593000/15,800,000

=10.08%(Approx).